# Flipped rotating axion non-minimally coupled to gravity: Baryogenesis and Dark Matter

### v1

@@ III Baryogenesis
The rotating axion can play role in generating the baryon asymmetry of the Universe Cohen and Kaplan (1987, 1988); Domcke et al. (2020). The important feature here is that the required rotation can be generated from the effective potential flipping direction after inflation, rather than explicit $U(1)$ breaking operators in the scalar potential with high radial field values as required in the usual Affleck-Dine scenarios Affleck and Dine (1985). Because of this tilt, the axion can attain a non-zero $\dot{\theta}$ and also climb up the potential barrier. The maximum kinetic energy after the end of inflation and flipping of the axion vacuum manifold, at a time say $t_{\rm m}$ can be written as
$\frac{1}{2}f^{2}\dot{\theta}_{\rm m}^{2}$ $\simeq 6\xi m_{P}^{2}H_{\rm end}^{2}$ (18)
$\implies\dot{\theta}_{\rm m}$ $\simeq\frac{\sqrt{12\xi}m_{P}H_{\rm end}}{f}\leavevmode\nobreak\ .$ (19)
Here, $\theta=\phi/f$ and $H_{\rm end}$ denotes the Hubble scale at the end of inflation ($H_{\rm end}\simeq 10^{13}\,$eV for GUT-scale inflation). The average $\dot{\theta}$ evolves as $\dot{\theta}\propto a^{-3}$. The presence of this non-zero velocity spontaneously breaks CPT in the expanding Universe, which in the presence of baryon number violating interactions in equilibrium can lead to the generation of baryon asymmetry, through spontaneous baryogenesis [^1] Here, we don’t consider transfer of asymmetry from decay of the rotating condensate Harigaya (2019); Barrie et al. (2022a, b); Barrie and Han (2024), since we require it to behave as dark matter at late times. Cohen and Kaplan (1987, 1988); Domcke et al. (2020).
The basic idea of spontaneous baryogenesis is that a dynamical pseudo-scalar field (the axion in our case) can generate an external chemical potential for quarks and/or leptons. This can happen through derivative couplings of the axion with fermion currents of the form $x_{\psi}\partial_{\mu}\phi j^{\mu}_{\psi}/f$, where $\psi$ indicates a fermion with $B-L$ charge $x_{\psi}$ and $j^{\mu}_{\psi}=\bar{\psi}\gamma_{\mu}\psi$ [^2] The backreaction from such a term to the axion equation of motion is suppressed for large $f$ Domcke et al. (2020), which is of our interest (cf. Eq. (16)). The presence of this coupling with a non-zero $\dot{\theta}$, accompanied by a B (or L) violating interaction in thermal equilibrium, causes an energy shift in particles (anti-particles) proportional to $\dot{\theta}(-\dot{\theta})$, giving rise to equilibrium values of the baryon or lepton asymmetry as $n_{B}\sim n_{L}\sim\dot{\theta}T^{2}$. The essential part of spontaneous baryogenesis is to generate a non-zero $\dot{\theta}$ sufficiently large for generating the observed asymmetry.
@@
As mentioned already, during kination the energy density of the rotating axion decreases as $\rho_{\phi}\propto t^{-2}$, which is also true of the dominant background density $\rho\propto t^{-2}$. Thus, we expect $\rho_{\phi}/\rho=\,$constant during kination. As in Ref. Bettoni et al. (2022), we expect that these rotations are not halted by the potential hill of $V_{\rm eff}$ in Eq. (7), because the latter is diminishing in height as $||V_{\rm eff}^{\prime\prime}||\propto t^{-2}$.
@@
where we considered Eq. (8). This remains (approximately) constant until the moment of reheating, after which the energy density of the dominant radiation background decreases as $\rho\propto a^{-4}$. The axion energy density continues to be dominated by the kinetic energy density, even though the effective potential is no more (because $R=0$ during the radiation era), such that $\rho_{\phi}\propto a^{-6}$. Thus, during this epoch we have
[…]
@@
11 11 $\simeq$ $\left.\frac{\rho_{\phi}}{\rho}\right|_{\rm eq}=\left(\frac{a_{\rm eq}}{a_{\rm ufr}}\right)\left.\frac{\rho_{\phi}}{\rho}\right|_{\rm ufr}$
@@
Using that $t_{\rm eq}\simeq 3.14\times 10^{35}\,$eV-1, we obtain [^4] Note that a large value of $f$, needed for the backreaction to be negligible, also leads to a small value of $M$.
@@
Let us consider some specific examples. Assume that the axion decay constant is at the scale of grand unification (GUT-scale), such that $f/m_{P}\sim 10^{-2}$. Assume also that inflation takes place at GUT-scale, such that $H\sim 10^{-5}\,m_{P}$. Then, Eq. (9) suggests that $110^{-4}\ll\xi<$ 1. Eq.  (33) and (31) give $M\sim 10^{-6}\,$eV and $m_{\phi}\sim 10^{-20}$ eV. Taking $\xi\sim 10^{-4}$, the bound in Eq. (34) suggests that $T_{\rm reh}>10^{-5}\,$eV. The latter is well satisfied, because the reheating temperature must be higher than the temperature at the time of BBN, $T_{\rm reh}>10^{-3}\,$eV.
Similarly, if we want to have $M\sim 1\,$TeV, then Eq. (33) suggests that $f\sim 10^{-8}m_{P}$, which corrresponds to $m_{\phi}\sim 40$ keV. The range in Eq. (9) becomes $10^{-16}<\xi\ll 1$. We select $\xi\sim 10^{-16}$. Now, Eq. (58) suggests that $H<10^{-7}m_{P}$ (see Sec. VIII). We choose $H\sim 10^{-8}m_{P}$, which means that the inflation energy scale is $V_{\rm end}^{1/4}\sim 10^{14}\,$eV. Then the bound in Eq. (34) becomes $T_{\rm reh}\gtrsim 10^{7}\,$eV. This agrees with the requirement in Eq. (47), which is discussed in the next section, that the peak of GW energy spectrum (cf. Eq. (V)) corresponding to kination does not disturb BBN, see Fig. 3.555Actually, the bound in Eq. (47) corresponds to GUT-scale inflation. It can be easily shown that when $V_{\rm end}^{1/4}\sim 10^{14}\,$eV the bound on $T_{\rm reh}$ is relaxed by about an order of magnitude. Thus we see that our model can accommodate $M\sim 1\,$TeV, which is a rather natural value at the electroweak energy scale. In terms of baryogenesis, choosing the extreme values of $\xi$ as in the above two examples, producing the observed baryon asymmetry requires (see Eq. (26)) $T_{\rm B-L}\sim 8\times 10^{7}$ GeV, for $T_{\rm reh}\sim 10^{7}$ GeV. Later in Sec. VI, we show an explicit model with such order of $T_{\rm B-L}$.
@@
If the bound in Eq. (34) is violated though, it does not mean that our mechanism for axion dark matter does not work. It simply means that the field never freezes. This can happen if $m_{\phi}$ catches up with Hubble ($H<m_{\phi}\sim M^{2}/f$) before the kinetic energy reaches the potential barrier $M^{4}$, thus switching to the vacuum potential in Eq. (3). In this case, while it is still rotating, it becomes subject to the potential in Eq. (3), stops rotating and begins coherent oscillations around its minimum at $\phi=0$. We expect this to occur if the reheating temperature is lower than $T_{\rm reh}^{\rm min}$ in Eq. (34).
@@ V Gravitational Waves
It is well known that the observed energy spectrum of primordial gravitational waves (GW) generated during inflation is affected by the thermal history after inflation Giovannini (1998); Gouttenoire et al. (2021a); Haque et al. (2021); Figueroa and Tanin (2019); Chen et al. (2024). The GW amplitudes are damped on subhorizon scales, i.e., $h_{k}(\tau)\propto k^{-3/2}a_{k}/a(\tau)$, where $k^{-3/2}$ is due to the fact that the GW power spectrum is nearly scale-invariant in the vanilla slow-roll inflation (as in our model). Here the subscript ‘$k$’ denotes the moment that the length scale of wavenumber $k$ reeneters the horizon during the hot Big Bang. Using the relationship [^6] Using $a\propto t^{2/3(1+w)}$ and $H\propto 1/t$. $k=a_{k}H_{k}\propto a^{-(1+3w)/2}$, namely, replacing $a_{k}$ by $k^{-2/(1+3w)}$, we derive the GW energy spectrum as [^7] For the detailed calculations, readers may refer to, e.g., Ref. Chen et al. (2024).
@@
$\Omega_{\rm GW}(\tau_{0},\nu)\simeq\Omega_{\rm GW}^{\rm RD}\left\{\begin{matrix}&\nu/\nu_{\rm reh}\leavevmode\nobreak\ ,\quad&&\nu_{\rm reh}<\nu<\nu_{\rm end}\\ \\ &1\leavevmode\nobreak\ ,&&\nu_{\rm eq}<\nu<\nu_{\rm reh}\\ \\ &(\nu_{\rm eq}/\nu)^{2}\leavevmode\nobreak\ ,&&\nu_{0}<\nu<\nu_{\rm eq}\end{matrix}\right.\leavevmode\nobreak\ ,$ (36)
@@
where $\Omega_{\rm\gamma}h^{2}\simeq 2.47\times 10^{-5}$ is the relic density of the radiation measured today, and $\nu_{\rm BBN}\sim 10^{-11}{\rm Hz}$. Applying Eqs. (36) and (V), we preform the above frequency integral and obtain
@@
where we used Eq. (40). Using the current bound from Planck ($\Delta N_{\rm eff}=0.17$), we get the lower bound on the reheating temperature and the peak of GW spectrum (V) as
@@
$0.06$ CMB-S4 Abazajian et al. (2019),PICO Hanany et al. (2019) $4.8\times 10^{7}\,\mathrm{GeV}$
@@ VI A Concrete Example
Figure 4: Predictions for cogenesis of dark matter and baryon asymmetry in the Type I seesaw setup with Majoron, considering $M_{N}=10^{7},5\times 10^{9}$ GeV in the left and right panel, respectively. The gray-shaded regions are ruled out from the BBN bound on GW (see Eq. (48)), whereas in the green-shaded regions $\xi$ lies outside the allowed range given by Eq. (9). Above the blue-dashed line, the symmetry of the scalar potential might get restored (see discussion at the end of Sec. II). Below the red solid contour in the right panel plot, the B-L violating interactions are out of equilibrium because of faster expansion of the Universe (see text). The different black lines indicate the future sensitivity reaches of several experiments for $\Delta N_{\rm eff}$ with solid, dashed, dot-dashed and dotted corresponding to BBN+CMB, CMB-S4/PICO, CMB-HD, COrE/EUCLID respectively (see Table 1).
@@
where $N$ indicates the right-handed neutrinos and $\mathcal{H},l$ are the SM Higgs and lepton doublets respectively. $i,j,\alpha$ run from 1 to 3. Note that the angular mode of the complex scalar $\Phi$ plays the role of Majoron, and undergoes the dynamics discussed in the preceding sections. The Majorana masses of $N$ are generated from $\Phi$ as $M_{N}=Y_{N}f/\sqrt{2}$, giving light neutrino mass $m_{\nu}$ through the usual seesaw mechanism as $m_{\nu}\sim Y_{D}^{2}\leavevmode\nobreak\ v^{2}/M_{N}$, where $v$ is the vacuum expectation value (VEV) of the standard model Higgs. Note that the light-neutrino mass is related to the Majoron mass as $m_{\nu}\sim 3\times 10^{4}\leavevmode\nobreak\ \frac{Y_{D}^{2}}{Y_{N}}\left(\frac{m_{\phi}}{\text{GeV}}\right)^{1/4}\left(\frac{\text{GeV}}{m_{P}}\right)^{3/4}\frac{v^{2}}{\text{GeV}}$, where we used condition for the dark matter abundance (Eq. (33)).
@@
In Fig. 4, we show the viable parameter space for generating the observed DM relic along with the baryon asymmetry in the $m_{\phi}$-$T_{\rm reh}$ plane, considering two benchmark values of the RHN mass: $M_{N}=10^{7}$ GeV, $5\times 10^{9}$ GeV in the left and right panels, respectively. For these values of mass, considering $m_{\nu}\sim 0.05$ eV, the Dirac Yukawa couplings are fixed through seesaw as $Y_{D}\sim 7\times 10^{-5}$ and $10^{-3}$, respectively. For a given value of the Majoron mass, the decay constant is fixed through Eq. (33), which is denoted in the upper X-axes. The gray shaded regions are ruled out from overproduction of GW, as given by Eq. (48). The red contours indicate different values of $\xi$ along which we can have cogenesis of DM and the baryon asymmetry, following Eqs. (24) and (33).
@@
For sufficiently low reheating temperatures, the Hubble rate turns large such that the inverse decays cannot be in equilibrium. The region below the red thick contour in the right panel plot for $M_{N}=5\times 10^{9}$ GeV indicates this, while in the left panel the bound from GW is much stronger. In the green-shaded regions $\xi$ lies outside the allowed range given by Eqs. (9) and (26). The lower bound on $\xi$ becomes important for a higher value of decoupling temperature (cf. Eq. (26)), as seen in the right panel plot for the heavier RHN mass. Above the blue-dashed contour, the value of $f$ becomes lower than the maximum temperature of the bath $T_{\rm max}$, which can potentially disrupt the VEV of the scalar potential as discussed earlier. As we will see in the next two sections, it is favorable for $\xi$ to lie close to the lower bound of the range in Eq. (9). Thus, the benchmark chosen in the left-panel ($M_{N}=10^{7}$ GeV) becomes unfavorable due to constraints from GW overproduction.
@@
The constraints discussed above are model-independent, once the decoupling temperature $T_{\rm B-L}$ is given, which in our case is primarily determined by the RHN mass scale $M_{N}$. Apart from these constraints, in order to have the Majoron as the dark matter candidate in our model, it needs to be cosmologically stable. The Majoron can decay into light neutrinos with a decay rate given by $\Gamma_{\phi\rightarrow\nu\nu}\simeq\frac{m_{\phi}}{16\pi f^{2}}\sum_{j}m_{\nu_{j}}^{2}$. Thus, the stability criteria is given by $\Gamma_{\phi\rightarrow\nu\nu}^{-1}\gtrsim\tau_{U}$, where $\tau_{U}\simeq 250$ GYr indicates the age of the Universe. Considering the current upper bound on the sum of neutrino mass $\sum_{j}m_{\nu_{j}}^{2}\simeq(0.05\text{eV})^{2}$, we get the upper bound on the Majoron mass as $m_{\phi}\lesssim 30$ MeV. Since, the interactions of Majoron with the standard model originate from the $Y_{D}$ term, the interactions are suppressed by the light neutrino mass $m_{\nu}$.
@@
Hence, testing our above predictions of light Majoron mass of sub-eVs becomes highly challenging. However, as discussed in Section V, our scenario can be probed through the blue-tilted gravitational wave spectra appearing because of the kination domination era. The requirement of cogenesis of baryon asymmetry and dark matter abundance predicts specific values of the reheating temperature (see Fig. 4), which determines the frequency $\nu_{\rm reh}$ given by Eq. (40), which in turn indicates the start of the blue-tilted spectra. Although tightly constrained by BBN bounds as discussed earlier, such a positive slope can be tested Chen et al. (2024) by future experiments like BBO Crowder and Cornish (2005); Corbin and Cornish (2006), UDECIGO Sato et al. (2017); Ishikawa et al. (2021), and also by proposed high-frequency resonant-cavity experiments Ringwald et al. (2021); Ringwald and Tamarit (2022) in the MHz-GHz regime, see Ref. Aggarwal et al. (2021) for a review.
VII Fragmentation of axion

### v2

@@ III Baryogenesis
The rotating axion can play a role in generating the baryon asymmetry of the Universe Cohen and Kaplan (1987, 1988); Domcke et al. (2020). The important feature here is that the required rotation can be generated from the effective potential flipping direction after inflation, rather than explicit $U(1)$ breaking operators in the scalar potential with high radial field values as required in the usual Affleck-Dine scenarios Affleck and Dine (1985). Because of this tilt, the axion can attain a non-zero $\dot{\theta}$ and also climb up the potential barrier, as we saw earlier (cf. Fig. 2). The amplitude of $\dot{\theta}$ scales as $a^{-3}$, i.e. $\dot{\theta}(t)||=\dot{\theta}_{m}\left(a_{m}/a(t)\right)^{3}$. Recall that in the limit $\xi\gg\left(f/m_{P}\right)^{2}$, $\dot{\theta}_{m}$ is given by Eq. (21). The presence of this non-zero velocity spontaneously breaks CPT in the expanding Universe, which in the presence of baryon number violating interactions in equilibrium can lead to the generation of baryon asymmetry, through spontaneous baryogenesis [^4] Here, we don’t consider transfer of asymmetry from decay of the rotating condensate Harigaya (2019); Barrie et al. (2022a, b); Barrie and Han (2024), since we require it to behave as dark matter at late times. Cohen and Kaplan (1987, 1988); Domcke et al. (2020).
The basic idea of spontaneous baryogenesis is that a dynamical pseudo-scalar field (the axion in our case) can generate an external chemical potential for quarks and/or leptons. This can happen through derivative couplings of the axion with fermion currents of the form $x_{\psi}\partial_{\mu}\phi j^{\mu}_{\psi}/f$, where $\psi$ indicates a fermion with BLB-L charge $x_{\psi}$ and $j^{\mu}_{\psi}=\bar{\psi}\gamma_{\mu}\psi$. [^5] The backreaction from such a term to the axion equation of motion is suppressed for large $f$ Domcke et al. (2020), which is of our interest (cf. Eq. (16)). The presence of this coupling with a non-zero $\dot{\theta}$, accompanied by a B (or L) violating interaction in thermal equilibrium, causes an energy shift in particles (anti-particles) proportional to $\dot{\theta}(-\dot{\theta})$, giving rise to equilibrium values of the baryon or lepton asymmetry as $n_{B}\sim n_{L}\sim\dot{\theta}T^{2}$. The essential part of spontaneous baryogenesis is to generate a non-zero $\dot{\theta}$ sufficiently large for generating the observed asymmetry.
@@
As mentioned already, during kination the energy density of the rotating axion decreases as $\rho_{\phi}\propto t^{-2}$, which is also true of the dominant background density $\rho\propto t^{-2}$. Thus, we expect $\rho_{\phi}/\rho=\,$constant during kination. As in Ref. Bettoni et al. (2022), we expect that these rotations are not halted by the potential hill of $V_{\rm eff}$ in Eq. (7), because the latter is diminishing in height as $V_{\rm eff}^{\prime\prime}||\propto t^{-2}$.
@@
where we considered Eq. (8). This remains (approximately) constant until the moment of reheating, after which the energy density of the dominant radiation background decreases as $\rho\propto a^{-4}$. The axion energy density continues to be dominated by the kinetic energy density, even though the effective potential is no more (because $R=0$ during the radiation era [^6] To be precise, quantum effects arising because of the trace anomaly of non-Abelian gauge theories Kajantie et al. (2003) induce a non-zero $R$, which for the SM can be approximated as Davoudiasl et al. (2004); Takahashi and Yamada (2015); Saikawa and Shirai (2018); Berbig (2024b) $R\simeq\frac{486}{19\pi^{2}}\alpha_{s}(T)^{2}H^{2}\,\simeq 0.037\,H^{2}$, where $\alpha_{s}(T\,\sim{\rm GeV}){\,\simeq 0.12}$ is the strong coupling constant Huston et al. (2023); Deur et al. (2022). The non-minimal coupling introduces a mass-squared of order $\sim\xi(m_{P}/f)^{2}R$ (cf. Eq. (4)). For $\xi\sim(f/m_{P})^{2}$, as suggested by Eq. (69), we find that the trace anomaly gives rise to an effective mass ${\cal O}(10^{-1})H$. Being smaller than $H$, such mass does not affect the dynamics of the axion, namely the freezing at $H_{\rm fr}$ (cf. Eq. (32)) and thawing when $H_{\rm ufr}\simeq M^{2}/f$ (cf. Eq. (33)). ), such that $\rho_{\phi}\propto a^{-6}$. Thus, during this epoch we have
$\frac{\rho_{\phi}}{\rho}=\left(\frac{a_{\rm reh}}{a}\right)^{2}\left.\frac{\rho_{\phi}}{\rho}\right|_{\rm reh}=\frac{H(t)}{H_{\rm reh}}\,2\xi\,$, (31)
@@
$1$ $\simeq$ $\left.\frac{\rho_{\phi}}{\rho}\right|_{\rm eq}=\left(\frac{a_{\rm eq}}{a_{\rm ufr}}\right)\left.\frac{\rho_{\phi}}{\rho}\right|_{\rm ufr}$
@@
Using that $t_{\rm eq}\simeq 3.14\times 10^{35}\,$GeV-1, we obtain [^8] Note that a large value of $f$, needed for the backreaction to be negligible, also leads to a small value of $M$.
@@
Let us consider a specific example. Assume that the axion decay constant is at the scale of grand unification (GUT-scale), such that $f/m_{P}\sim 10^{-2}$. Assume also that inflation takes place at GUT-scale, such that $H\sim 10^{-5}\,m_{P}$. Then, Eq. (9) suggests that $10^{-4}\ll\xi<1$. Eqs. (33) and (35) give $M\sim 10^{-6}\,$GeV and $m_{\phi}\sim 10^{-20}\,$eV. Taking $\xi\sim 10^{-4}$, the bound in Eq. (36) suggests that $T_{\rm reh}>10^{-5}\,$GeV. The latter is well satisfied, because the reheating temperature must be higher than the temperature at the time of BBN, $T_{\rm reh}>10^{-3}\,$GeV. In terms of baryogenesis, when choosing the extreme value of $\xi$ as in the above example, producing the observed baryon asymmetry requires (see Eq. (28)) $T_{\rm B-L}\simeq 8\times 10^{7}$ GeV, for $T_{\rm reh}\sim 10^{7}$ GeV. Later in Sec. VI, we show an explicit model with such order of $T_{\rm B-L}$.
@@
If the bound in Eq. (36) is violated though, it does not mean that our mechanism for axion dark matter does not work. It simply means that the field never freezes. This can happen if $m_{\phi}$ catches up with Hubble ($H<m_{\phi}\sim M^{2}/f$) before the kinetic energy reaches the potential barrier $M^{4}$, thus switching to the vacuum potential in Eq. (3). In this case, while it is still rotating, the axion becomes subject to the potential in Eq. (3), stops rotating and begins coherent oscillations around its minimum at $\phi=0$. We expect this to occur if the reheating temperature is lower than $T_{\rm reh}^{\rm min}$ in Eq. (36).
@@ V Gravitational Waves
It is well known that the observed energy spectrum of primordial gravitational waves (GW) generated during inflation is affected by the thermal history after inflation Giovannini (1998); Gouttenoire et al. (2021a); Haque et al. (2021); Figueroa and Tanin (2019); Chen et al. (2024). The GW amplitudes are damped on subhorizon scales, i.e., $h_{k}(\tau)\propto k^{-3/2}a_{k}/a(\tau)$, where $k^{-3/2}$ is due to the fact that the GW power spectrum is nearly scale-invariant in the vanilla slow-roll inflation (as in our model). Here the subscript ‘$k$’ denotes the moment that the length scale of wavenumber $k$ reeneters the horizon during the hot Big Bang. Using the relationship [^9] Obtained using $a\propto t^{2/3(1+w)}$ and $H\propto 1/t$. $k=a_{k}H_{k}\propto a^{-(1+3w)/2}$, namely, replacing $a_{k}$ by $k^{-2/(1+3w)}$, we derive the GW energy spectrum as [^10] For the detailed calculations, readers may refer to, e.g., Ref. Chen et al. (2024).
@@
[…]
@@
where $\Omega_{\rm\gamma}h^{2}\simeq 2.47\times 10^{-5}$ corresponds to the relic density of the radiation measured today, and $\nu_{\rm BBN}\sim 10^{-11}{\rm Hz}$. Applying Eqs. (38) and (V), we obtain
@@
where we used Eq. (42). Using the current value from the Planck observations ($\Delta N_{\rm eff}=0.17$), we get the lower bound on the reheating temperature and the peak of GW spectrum (V) as
@@
$0.06$ CMB-S4 Abazajian et al. (2019),PICO Hanany et al. (2019), CMB-Bharat  Adak et al. (2022) $4.8\times 10^{7}\,\mathrm{GeV}$
@@ VI A Concrete Example
Figure 5: Predictions for cogenesis of dark matter and baryon asymmetry in the Type I seesaw setup with Majoron, considering $M_{N}=10^{7},5\times 10^{9}$ GeV in the left and right panel, respectively. The gray-shaded regions are ruled out from the BBN bound on GW (see Eq. (50)), whereas in the green-shaded regions $\xi$ lies outside the allowed range given by Eq. (9). Above the blue-dashed line, the symmetry of the scalar potential might get restored (see discussion at the end of Sec. II). Below the red solid contour in the right panel plot, the B-L violating interactions are out of equilibrium because of faster expansion of the Universe (see text). The different black lines indicate the future sensitivity reaches of several experiments for $\Delta N_{\rm eff}$ with solid, dashed, dot-dashed and dotted corresponding to BBN+CMB, CMB-S4/PICO/CMB-Bharat, CMB-HD, COrE/EUCLID respectively (see Table 1).
@@
where $N$ indicates the right-handed neutrinos and $\mathcal{H},l$ are the SM Higgs and lepton doublets respectively. $i,j,\alpha$ run from 1 to 3. Note that the angular mode of the complex scalar $\Phi$ plays the role of Majoron, and undergoes the dynamics discussed in the preceding sections. The Majorana masses of $N$ are generated from $\Phi$ as $M_{N}=Y_{N}f/\sqrt{2}$, giving light neutrino mass $m_{\nu}$ through the usual seesaw mechanism as $m_{\nu}\sim Y_{D}^{2}\leavevmode\nobreak\ v^{2}/M_{N}$, where $v$ is the vacuum expectation value (VEV) of the standard model Higgs.
@@
In Fig. 5, we show the viable parameter space for generating the observed DM relic along with the baryon asymmetry in the $m_{\phi}-T_{\rm reh}$ plane, considering two benchmark values of the RHN mass: $M_{N}=10^{7}$ GeV, and $5\times 10^{9}$ GeV in the left and right panels, respectively. For these values of mass, considering $m_{\nu}\sim 0.05$ eV, the Dirac Yukawa couplings are fixed through seesaw as $Y_{D}\sim 7\times 10^{-5}$ and $10^{-3}$, respectively. For a given value of the Majoron mass, the decay constant is fixed through Eq. (35), which is denoted in the upper X-axes. The gray shaded regions are ruled out from overproduction of GWs, as given by Eq. (50). The red contours indicate different values of $\xi$ along which we can have cogenesis of DM and the baryon asymmetry, following Eqs. (26) and (35).
@@
For sufficiently low reheating temperatures, the Hubble rate turns large such that the inverse decays cannot be in equilibrium. The region below the red thick contour in the right panel plot for $M_{N}=5\times 10^{9}$ GeV indicates this, while in the left panel the bound from GW overproduction is much stronger. In the green-shaded regions $\xi$ lies outside the allowed range given by Eqs. (9) and (28). The lower bound on $\xi$ becomes important for a higher value of decoupling temperature (cf. Eq. (28)), as seen in the right panel plot for the heavier RHN mass. Above the blue-dashed contour, the value of $f$ becomes lower than the maximum temperature of the bath $T_{\rm max}$, which can potentially disrupt the VEV of the scalar potential as discussed earlier. As we will see in the next two sections, it is favorable for $\xi$ to lie close to the lower bound of the range in Eq. (9). Thus, the benchmark chosen in the left-panel ($M_{N}=10^{7}$ GeV) becomes unfavorable due to constraints from GW overproduction.
@@
The constraints discussed above are model-independent, once the decoupling temperature $T_{\rm B-L}$ is given, which in our case is primarily determined by the RHN mass scale $M_{N}$. Apart from these constraints, in order to have the Majoron as the dark matter candidate in our model, it needs to be cosmologically stable. The Majoron can decay into light neutrinos with a decay rate given by $\Gamma_{\phi\rightarrow\nu\nu}\simeq\frac{m_{\phi}}{16\pi f^{2}}\sum_{j}m_{\nu_{j}}^{2}$. Thus, the stability criteria is given by $\Gamma_{\phi\rightarrow\nu\nu}^{-1}\gtrsim\tau_{U}$, where $\tau_{U}\simeq 250$ Gyr Audren et al. (2014); Enqvist et al. (2020); Nygaard et al. (2021); Alvi et al. (2022); Simon et al. (2022). Considering the current upper bound on the sum of neutrino mass $\sum_{j}m_{\nu_{j}}^{2}\simeq(0.05\text{eV})^{2}$, we get the upper bound on the Majoron mass as $m_{\phi}\lesssim 30$ MeV. Since, the interactions of Majoron with the standard model originate from the $Y_{D}$ term, the interactions are suppressed by the light neutrino mass $m_{\nu}$.
@@
Hence, testing our above predictions of light Majoron mass of sub-eVs becomes highly challenging. However, as discussed in Section V, our scenario can be probed through the blue-tilted gravitational wave spectra appearing because of the kination domination era. The requirement of cogenesis of baryon asymmetry and dark matter abundance predicts specific values of the reheating temperature (see Fig. 5), which determines the frequency $\nu_{\rm reh}$ given by Eq. (42), which in turn indicates the start of the blue-tilted spectra. Although tightly constrained by BBN bounds as discussed earlier, such a positive slope can be tested Chen et al. (2024) by future experiments like BBO Crowder and Cornish (2005); Corbin and Cornish (2006), UDECIGO Sato et al. (2017); Ishikawa et al. (2021), and also by proposed high-frequency resonant-cavity experiments Ringwald et al. (2021); Ringwald and Tamarit (2022) in the MHz-GHz regime, see Ref. Aggarwal et al. (2021) for a review. Moreover, the $\Delta N_{\rm eff}$ values from the GW background, can be within reach of several near-future CMB experiments (cf. Table 1), which are shown by different black lines in Fig. 5.
Finally, our example also has some predictions for light neutrino mass as follows. Using the seesaw formula ($m_{\nu}\sim Y_{D}^{2}\leavevmode\nobreak\ v^{2}/M_{N}$) and the condition for the dark matter abundance (cf. Eq. (35)), the light-neutrino mass can be related to the Majoron mass as
$m_{\nu}\sim 3\times 10^{4}\,v\,\frac{Y_{D}^{2}}{Y_{N}}\left(\frac{m_{\phi}}{\text{GeV}}\right)^{1/4}\left(\frac{\text{GeV}}{m_{P}}\right)^{3/4}\frac{v}{\text{GeV}}$ (53)
From the experimental constraints of neutrinoless double $\beta$-decay from KATRIN, the direct neutrino mass measurement gives Aker et al. (2022), $m_{\nu}\leq 0.8$ eV and future sensitivity can reach upto $m_{\nu}\leq 0.2$ eV. Thus, Majoron masses below $\mathcal{O}$ (GeV), can be indirectly tested at these neutrino experiments. The scenario becomes more predictive considering that it gives rise to the baryon asymmetry during the rotation phase, since $Y_{N}$ is determined ($T_{\rm B-L}\sim\frac{M_{N}}{z_{\rm fo}}\sim\frac{Y_{N}f}{z_{\rm fo}}$) imposing Eq. (26). Considering $\xi\sim(f/m_{P})^{2}$ (cf. Eq. (69)), the light neutrino mass is predicted to be
$\hskip-11.38092ptm_{\nu}\sim c_{B}^{1/2}\left(\frac{v}{246\leavevmode\nobreak\ \text{GeV}}\right)\left(\frac{10}{z_{\rm fo}}\right)\left(\frac{10^{7}\leavevmode\nobreak\ \text{GeV}}{T_{\rm reh}}\right)^{1/2}$
$\times\left(\frac{Y_{D}}{6.3\times 10^{-2}}\right)^{2}\text{eV}.$ (54)
Hence, the setup has complementary tests in GW, CMB and neutrino experiments.
VII Axion Fragmentation